OpenStudy (anonymous):

*calculus 3* Let vector U = 4i - 2k and let vector V = i-2j+k _Find the magnitude (norm) of 3U - 2V - aka II 3U - 2V II _Two vectors of length 2 which are also parallel to 3U - 2V _The components and projections of 3U along 2V

OpenStudy (michele_laino):

hint: \[3U - 2V = \left( {\begin{array}{*{20}{c}} {12}&0&{ - 6} \end{array}} \right) - \left( {\begin{array}{*{20}{c}} 2&{ - 4}&2 \end{array}} \right)\]

OpenStudy (anonymous):

I understand that part. i calculated the magnitude and got 180^1/2 However, i am not sure how to solve the second or third part.

OpenStudy (michele_laino):

We have: \[\begin{gathered} W = 3U - 2V = \left( {\begin{array}{*{20}{c}} {12}&0&{ - 6} \end{array}} \right) - \left( {\begin{array}{*{20}{c}} 2&{ - 4}&2 \end{array}} \right) \hfill \\ W = \left( {\begin{array}{*{20}{c}} {10}&4&{ - 8} \end{array}} \right) \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

a vector parallel to vector W is the subsequent vector: \[kW = \left( {\begin{array}{*{20}{c}} {10k,}&{4k,}&{ - 8k} \end{array}} \right)\] so we have to determine the value of k such that the length of kW is equal to 2

OpenStudy (anonymous):

Would k= (2/(180^1/2))? I first made the magnitude equal to 1 by dividing it by 180^1/2, then I multiplied by 2 to get me a vector the length of 2.

OpenStudy (michele_laino):

we have 2 values of k since we can write: \[\left\| W \right\| = \left| k \right|\sqrt {180} \] so: \[k = \pm \frac{2}{{\sqrt {180} }}\]

OpenStudy (michele_laino):

oops.. \[\left\| {kW} \right\| = \left| k \right|\sqrt {180} \]

OpenStudy (michele_laino):

In order to get the projection and component we have to compute the Fourier's coefficient of one vector with respect to the other

OpenStudy (michele_laino):

for example the projection of the vector 3U along the vector 2V is given by the subsequent formula: \[c2V = \frac{{\left\langle {3U,2V} \right\rangle }}{{\left\langle {2V,2V} \right\rangle }}2V\]

OpenStudy (michele_laino):

where: \[\left\langle { \cdot , \cdot } \right\rangle \] is the scalar product and c is the Fourier's coefficient: \[c = \frac{{\left\langle {3U,2V} \right\rangle }}{{\left\langle {2V,2V} \right\rangle }}\]

OpenStudy (michele_laino):

so we have to compute these quantities: \[{\left\langle {3U,2V} \right\rangle }\] and \[{\left\langle {2V,2V} \right\rangle }\]

OpenStudy (anonymous):

So I would be doing the dot product for the numerator (in this case 3U and 2V) and for the denominator (in this case 2V and 2V)?

OpenStudy (michele_laino):

yes! that's the coefficient c, then the projection is the vector: c 2V

OpenStudy (anonymous):

Ah ok, got it! Just hypothetically, if they wanted the reverse "find the projection of 2V along 3U - the Fourier coefficient would be \[c = \frac{ <2U, 3V> }{<3V,3V> }\]

OpenStudy (anonymous):

then I would multiply the coefficent by vector 3U

OpenStudy (anonymous):

oops i mixed up the variables for the fraction: 2V and 3U

OpenStudy (michele_laino):

the projection of the vector 2V along the vector 3U is: \[{c_1}3U = \frac{{\left\langle {2V,3U} \right\rangle }}{{\left\langle {3U,3U} \right\rangle }}3U\]

OpenStudy (anonymous):

Ok good! Thanks for the help! :)

OpenStudy (michele_laino):

we have to swap 2V with 3U

OpenStudy (michele_laino):

Thanks! :)